a)     Suppose that UNZA is a country with a population of 321,166,640 people of which the number of unemployed people is 121,166,640. The employment rate in this economy is 0.104, and the labour force participation rate 72.5%.


1
Expert's answer
2021-07-08T15:13:00-0400

(i)

"labor \space force =\frac{unemployed}{unemployment \space rate}"


"=\frac{121,166,640}{0.104}×100\\=116,505,384,600"


(ii) unemployed people "=121,166,640"


(iii)

"working age =\space{employed}{labor \space force\space participation \space rate}"


"=\frac{11,638,521,800}{0.0725}\\=160,531,335,200"